zarkam21:

When a certain amount of violet light and yellow light shine on a metal surface, only violet light produces an electric current. Write 3 - 4 sentences describing how an increase in the intensity of each type of light will affect the amount of current produced. How does quantization help to explain these observations?

Vocaloid:

this phenomenon is called the photoelectric effect so there are two factors of light that affect the ejection of electrons from the metal: intensity and frequency frequency determines whether electrons will be ejected or not (only high enough frequency light ejects electrons; if the frequency is too low, then no electrons are ejected regardless of intensity) intensity determines the rate at which electrons will be ejected

Vocaloid:

with that being said: - how does changing the intensity of violet light change the current? - how does changing the intensity of yellow light change the current?

Vocaloid:

 notice how current increases as intensity increases, but only if the frequency is high enough

zarkam21:

- how does changing the intensity of violet light change the current? increases - how does changing the intensity of yellow light change the current? increses

Vocaloid:

hint: going back to the original problem "only violet light produces an electric current. "

zarkam21:

- how does changing the intensity of violet light change the current? increases - how does changing the intensity of yellow light change the current? Does not

Vocaloid:

good and as for the quantization part, the reason why only violet light produces current is that light is quantized, and only photons with the right amount of energy (the right quanta) can eject a particle

zarkam21:

1 attachment
Vocaloid:

uh it's kinda hard to describe the shape but you gotta draw the field lines pointing north to south like this 

zarkam21:

so all these lines?

Vocaloid:

yeah something like that, make sure the arrows are pointing in the right direction (north to south) then for the force, I guess you could say the force at any location between the two magnets follows the direction of the field line, north to south

Vocaloid:

anyway for the last part of the question you just need to decide which two electric charges produce similar field lines, ex. two positives or two negatives or one positive one negative

zarkam21:

one positive and one negative?

Vocaloid:

good, or to be more specific, one proton one electron; electron has to be on the left and proton on the right to produce the same field lines in the same direction

zarkam21:

1 attachment
Vocaloid:

huh. what exactly are 1 and 2 representing? I think(?) the resistors are still in series so total resistance is just the sum

zarkam21:

36

Vocaloid:

12 is the voltage so it cannot be added to the resistors

zarkam21:

24

Vocaloid:

good, that's it I believe (units are ohms)

zarkam21:

okay so using this same diagram : c. In the circuit diagram above, meters 1 and 2 are connected as shown. Write 2 - 3 sentences identifying each type of meter and how it is connected with the 8 Ω resistor. d. In the circuit diagram above, predict which resistors will stop working (if any) when the switch is opened. Write 2 - 3 sentences explaining your reasoning.

Vocaloid:

c: well there are two types of meters, an ammeter and voltmeter, any ideas which is which? as a hint the ammeter must be in series with the resistors and the voltmeter must be in parallel

zarkam21:

series so ammeter

Vocaloid:

is the ammeter meter 1 or meter 2?

zarkam21:

1

nuts:

meter 1 is in parallel with the resistor :(

zarkam21:

2?

Vocaloid:

yeah, the ammeter must be meter 2 meaning the voltmeter is meter 1

zarkam21:

okay so for part d

zarkam21:

no

zarkam21:

I Have @Vocaloid

Vocaloid:

uh so have you given any thought to it yet? there are only a few possibilities, either they both keep working, they both stop working, or only one keeps working

zarkam21:

one stops working?

nuts:

why do you say that?

zarkam21:

actually all of them

zarkam21:

they all stop working because they are in a series?

nuts:

that's correct. Because the resistors are in series if either one of them fails, there is no closed path for current to go through any of them, so both stop working.

zarkam21:

A light wave moves through glass (n = 1.5) at an angle of 15°. What angle will it have when it moves from the glass into air (n = 1)?

nuts:

Do you remember snell's law?

zarkam21:

No

zarkam21:

is it the n1 equation. I think ?

nuts:

yeah. so \(\Huge \frac{\sin \theta_1}{\sin \theta_2}=\frac{n_2}{n_1}\). Well actually i don't know the critical angle but i'm pretty sure we use snell's law cause pretty sure refraction and not reflection occur.

zarkam21:

sin15/sin(theta)=1/1.5 theta=1.03

nuts:

your first equation is correct, but i'm not sure how you're getting theta=1.03

zarkam21:

oh 0.67?

nuts:

could you show your work?

zarkam21:

0.39?

zarkam21:

I think I got it now

Vocaloid:

might be a degrees/radians issue?

nuts:

Yeah it is. You've given the correct number in radians but you need to convert to degrees or use the correct mode on whatever calculator you're using

zarkam21:

22.35

nuts:

yup you're well on your way to design a spectrophotometer, all you need is to learn about dispersion now :)